# Privacy-Aware Sequential Learning

### v4

@@ 4.1 Main Results for Continuous Signals
Theorems 4.17 and 4.18 together demonstrate that, even in the presence of heterogeneous privacy budgets, the learning process remains efficient. Specifically, agents are expected to take a correct action in finite time, and the overall number of incorrect actions remains bounded. These results further confirm that smooth randomized response enables timely and reliable decision-making under privacy constraints.
Simulation under Heterogeneous Privacy Budgets
While our theory shows that stronger privacy concerns yield faster asymptotic learning (as $n\to\infty$), we ask how learning behaves in finite populations. Following Alaggan et al. (2015), Acharya et al. (2025), we classify agents by privacy: conservatives ($\varepsilon=0.1$), pragmatists ($\varepsilon=0.5$), and liberals ($\varepsilon=1$). We study five regimes: three homogeneous settings with fixed $\varepsilon\in{0.1,0.5,1}$, a heterogeneous setting with $\varepsilon\sim U(0,1)$, and a non-private baseline. For each regime we run five independent simulations; signals are normal with standard deviation $\sigma=3$. This design ties behavioral types to budgets and tests whether asymptotic predictions already manifest in finite samples.
Figure 5: Log-likelihood ratio dynamics under five privacy regimes. Homogeneous settings: conservatives $(\varepsilon=0.1)$, pragmatists $(\varepsilon=0.5)$, and liberals $(\varepsilon=1)$; heterogeneous setting: $\varepsilon\sim U(0,1)$; and a non-private baseline. Signals are normally distributed with $\sigma=3$.
Figure 5 illustrates the trajectory of the log-likelihood ratio $l_{n}$ as the number of agents increases. We observe that, consistent with the theoretical result, the log-likelihood ratio grows more rapidly under stronger privacy concerns (lower budgets). Intuitively, when privacy concern is stronger, agents introduce more randomization in their actions. Under the false state, this randomization causes agents to more frequently deviate from the majority action. Such deviations asymmetrically amplify the log-likelihood ratio between states, thereby accelerating information aggregation.
@@ 5 Conclusions
For continuous signals, privacy-aware agents adopt the smooth randomized response strategy, which adjusts noise based on the signal’s distance from the decision threshold. This ensures metric differential privacy while reducing unnecessary noise for signals and amplifying differences between actions. The strategy introduces asymmetry that makes correct actions more informative, enhancing belief updates and accelerating learning. As a result, asymptotic learning is not only preserved under privacy but also occurs at an improved rate of $\Theta_{\varepsilon}(\log n)$, compared to $\Theta(\sqrt{\log n})$ without privacy. These findings highlight how adaptive noise strategies can improve learning efficiency even under stringent privacy constraints.
@@
All R scripts used to generate the figures and reproduce the simulation results in this paper are openly available at our GitHub repository: https://github.com/YuxinLiu1997/Privacy-Preserving-Sequential-Learning.
@@ 6 Additional Related Work
$k=\lfloor\log_{\frac{u(1-p)+p(1-u)}{(1-u)(1-p)+up}}\frac{p}{(1-p)}\rfloor+1.$ (14)
@@
$l_{3}=\log\frac{\mathbb{P}(x_{1}=+1|\theta=+1)}{\mathbb{P}(x_{1}=+1|\theta=-1)}\frac{\mathbb{P}(x_{2}=+1|\theta=+1,x_{1}=+1)}{\mathbb{P}(x_{2}=+1|\theta=-1,x_{1}=+1)}=\log\left[\frac{u(1-p)+p(1-u)}{(1-u)(1-p)+up}\right]^{2}.$
@@
$l_{3}=\log\frac{\mathbb{P}(x_{1}=-1|\theta=+1)}{\mathbb{P}(x_{1}=-1|\theta=-1)}\frac{\mathbb{P}(x_{2}=-1|\theta=+1,x_{1}=-1)}{\mathbb{P}(x_{2}=-1|\theta=+1,x_{1}=-1)}=\log\left[\frac{(1-u)(1-p)+up}{u(1-p)+p(1-u)}\right]^{2}.$
@@
$\log\left[\frac{(1-u)(1-p)+up}{u(1-p)+p(1-u)}\right]^{2}+\log\frac{p}{1-p}<0,\log\left[\frac{u(1-p)+p(1-u)}{(1-u)(1-p)+up}\right]^{2}+\log\frac{1-p}{p}>0.$
@@
In total, if $k$ is the threshold for the start of the information cascade, then $k=\lfloor\log_{\frac{u(1-p)+p(1-u)}{(1-u)(1-p)+up}}\frac{p}{(1-p)}\rfloor+1$.
@@
When $\tilde{u}\neq\frac{1}{2}$, the probability of a correct cascade is given by:
$\mathbb{P}(\text{correct cascade})=\frac{\rho^{k}-1}{\rho^{2k}-1}.$
where $\rho=\frac{1-\tilde{u}}{\tilde{u}}$ and $\tilde{u}=u(1-p)+p(1-u)$. When $\tilde{u}=\frac{1}{2}$, the information cascade does not occur.
@@
For the case $\tilde{u}=\frac{1}{2}$, it is trivial. However, for $\tilde{u}\neq\frac{1}{2}$, we can reframe the problem as a Markov chain featuring two absorbing states: $k$ and k-k. The transition probabilities are defined as follows: $\mathbb{P}(z_{n}=m+1|z_{n-1}=m)=\tilde{u}$ and $\mathbb{P}(z_{n}=m-1|z_{n-1}=m)=1-\tilde{u}$, where $z_{n-1}$ is the difference between the occurrences of action +1+1 and action 1-1 in the observation history up to agent $n$. This formulation aligns with the classical gambler’s ruin problem.
@@
$s_{2k}(k)=\tilde{u}s_{2k}(k+1)+(1-\tilde{u})s_{2k}(k-1).$
@@
$s_{2k}(k)=\frac{\rho^{k}-1}{\rho^{2k}-1}.$
where $\rho=\frac{1-\tilde{u}}{\tilde{u}}$ and $\tilde{u}=u(1-p)+p(1-u)$. ∎
@@
where $\alpha=(1-p)/p$. Thus, $I_{k}=(\varepsilon_{k+1},\varepsilon_{k}]$, and $\varepsilon_{k}=\log\frac{1-v_{k}}{v_{k}}$. By lemma 8.2 take the first derivative of $s_{2k}(k)$ with respect to $\rho$ for $k\geq 2$ and $\rho\in(0,1)$:
$\frac{ds_{2k}(k)}{d\rho}=\frac{3k\rho^{3k-2}-k\rho^{k-1}-2k\rho^{2k-1}}{(\rho^{2k}-1)^{2}}<0.$
Next, we take the first derivative of $\rho$ and $u$ with respect to $u$ and $\varepsilon$, respectively. Then $\frac{d\rho}{du}>0$ and $\frac{du}{d\varepsilon}<0$. Consequently. $\frac{ds_{2k}(k)}{d\varepsilon}>0$, With an increase in the differential privacy budget, the probability of a correct cascade also increases.
9 Proof of Theorem 4.2: Randomized Response Strategy for the Continuous Model
we can derive the decision for agent $n$ before adding any noise. A simple calculation shows that agent $n$ takes action +1+1 iff $l_{n}+L_{n}>0$ which implies
@@
$L_{n}=\log{\frac{\mathbb{P}(\theta=+1|s_{n})}{\mathbb{P}(\theta=-1|s_{n})}}=\log\frac{f(s_{n}|\theta=+1)}{f(s_{n}|\theta=-1)}=\log{\frac{e^{-(s_{n}-1)^{2}/2}}{e^{-(s_{n}+1)^{2}/2}}}=e^{2s_{n}}.$
@@
where $\tilde{a_{n}}=-a_{n}$, then $u_{n}=\frac{1}{1+e^{\varepsilon}}$.
10 Proof of Theorem 4.3: Failure of Asymptotic Learning
Without loss of generality, we assume $\theta=-1$, then the likelihood ratio of the public belief
@@
$=(1-u)G_{-}(-l_{\infty})+u(1-G_{-}(-l_{\infty}))$
$=1-u.$
In all, because agents always randomize actions for the purpose of protecting privacy with constant probability, asymptotic learning will not occur.
@@
where $D_{+}(l_{n})=\log\frac{\mathbb{P}(x_{n}=+1|l_{n},\theta=+1)}{\mathbb{P}(x_{n}=+1|l_{n},\theta=-1)}$, $G_{-}(-l_{n})=a(e^{\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}-e^{-\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}})e^{-\varepsilon l_{n}}$, and $a=\frac{1}{2}$.
@@
$\mathbb{P}_{\pi,-}(\sigma<+\infty)\geq\frac{4\alpha}{1+2\alpha}.$
@@
$\frac{\phi(\pi|x)}{1-\phi(\pi|x)}=\frac{\pi}{1-\pi}\cdot\frac{\mathbb{P}(x_{n}=+1|\pi,\theta=+1)}{\mathbb{P}(x_{n}=+1|\pi,\theta=-1)}.$
@@
$\mathbb{P}(x_{n}=+1|\pi,\theta=+1)<\mathbb{P}(x_{n}=+1|\pi,\theta=-1),$
@@
By Lemma 14.1, $\pi_{n+1}^{*}\leq\frac{1}{2}$, and by Lemma 14.3, we have
@@ 16 The Evolution of Log-Likelihood Ratio
Proof of Theorems 4.17 and 4.18: Learning Efficiency—Finite Expected Time to First Correct Action and Finite Expected Total Number of Incorrect Actions with Heterogeneous Privacy Budgets
Theorems 4.17 and 4.18 mirror the structure and argument of Theorems 4.12 and 4.13, which analyze learning efficiency in the homogeneous privacy setting. The key steps in the proof remain the same: by substituting the asymptotic learning rate result in Theorem 4.15, we obtain the corresponding expression of the expectations. Therefore, we omit the full proof and show only that the expectations in both theorems are finite.
@@ 20 Extension to Pufferfish Privacy
$u_{n}(s_{n})=\frac{e^{-k\varepsilon}}{1+e^{\varepsilon}},\quad\text{if }s_{n}\in(t(l_{n})+ka,\ t(l_{n})+(k+1)a)\cup(t(l_{n})-ka,\ t(l_{n})-(k+1)a),\ k\in\mathbb{N}.$
@@
When the parameter $a$ increases, if the ratio $\frac{a}{\varepsilon}$ remains constant, then the asymptotic learning rate $f(n)=\frac{2a}{\varepsilon\sigma^{2}}\log(C_{\varepsilon,a}n)$ increases. This is because a larger $a$ induces a more asymmetric smooth randomized response, which improves the informativeness of the observed actions and thus accelerates the overall learning process. However, this improvement comes with a caveat: as $a\to\infty$, the term $C_{\varepsilon,a}$ tends to zero, which causes $f(n)$ to diverge. Therefore, we restrict our attention to the regime where $\frac{\varepsilon}{a}$ is constant and bounded away from zero to ensure that the learning rate remains well defined and informative.
@@
where $C_{1}$ is a positive constant that does not depend on $\varepsilon$ and $\zeta\left(x\right)=\sum_{n=1}^{\infty}n^{-x}$. If $\frac{2a}{\varepsilon\sigma^{2}}>1$, the series converges and thus the expected time to the first correct action is finite:
@@
where $C_{2}$ is a positive constant that does not depend on $\varepsilon$. If and only if $\frac{2a}{\varepsilon\sigma^{2}}>1$, the series converges and thus the expected total number of incorrect actions is finite:
@@
Although a higher value of $a$ makes it easier to ensure the finiteness of both $\mathbb{E}[\tau]$ and $\mathbb{E}[W]$, this benefit is not without potential drawbacks. In particular, as $a\to\infty$, the term $C_{\varepsilon,a}^{-\frac{2}{\varepsilon\sigma^{2}}}$ diverges to infinity. Therefore, our results are valid only under the condition that $\frac{\varepsilon}{a}$ is constant and bounded away from zero. This ensures that the learning efficiency bounds remain meaningful and finite.

### v5

@@ 4.1 Main Results for Continuous Signals
where $C_{2}$ and $\tilde{C}$ are positive constants. The series in Equation 13 is finite.
Theorems 4.15 and 4.16 together demonstrate that, even in the presence of heterogeneous privacy budgets, the learning process remains efficient. Specifically, agents are expected to take a correct action in finite time, and the overall number of incorrect actions remains bounded. These results further confirm that smooth randomized response enables timely and reliable decision-making under privacy constraints.
Simulation under Heterogeneous Privacy Budgets.
While our theory shows that stronger privacy concerns yield faster asymptotic learning (as $n\to\infty$), we ask how learning behaves in finite populations. Following Alaggan et al. (2015), Acharya et al. (2025), we classify agents by privacy: conservatives ($\varepsilon=0.1$), pragmatists ($\varepsilon=0.5$), and liberals ($\varepsilon=1$). We study five regimes: three homogeneous settings with fixed $\varepsilon\in{0.1,0.5,1}$, a heterogeneous setting with $\varepsilon\sim U(0,1)$, and a non-private baseline. For each regime we run five independent simulations; signals are normal with standard deviation $\sigma=1$. This design ties behavioral types to budgets and tests whether asymptotic predictions already manifest in finite samples.
Figure 5: Log-likelihood ratio dynamics under five privacy regimes. Homogeneous settings: conservatives $(\varepsilon=0.1)$, pragmatists $(\varepsilon=0.5)$, and liberals $(\varepsilon=1)$; heterogeneous setting: $\varepsilon\sim U(0,1)$; and a non-private baseline. Signals are normally distributed with $\sigma=1$.
Figure 5 illustrates the trajectory of the log-likelihood ratio $l_{n}$ as the number of agents increases. We observe that, consistent with the theoretical result, the log-likelihood ratio grows more rapidly under stronger privacy concerns (lower budgets). Intuitively, when privacy concern is stronger, agents introduce more randomization in their actions. This randomization increases the gap between action distributions under the two states: under the true state, agents conform to the majority action more reliably, whereas under the false state, their actions are noisier and deviate more often. This asymmetry amplifies the log-likelihood ratio between states, thereby accelerating information aggregation.
@@ 5 Conclusions
For continuous signals, privacy-aware agents adopt the smooth randomized response strategy, which adjusts noise based on the signal’s distance from the decision threshold. This ensures metric differential privacy while reducing unnecessary noise for signals and amplifying informativeness differences between actions. The strategy introduces asymmetry that makes correct actions more informative, enhancing belief updates and accelerating learning. As a result, asymptotic learning is not only preserved under privacy but also occurs at an improved rate of $\Theta_{\varepsilon}(\log n)$, compared to $\Theta(\sqrt{\log n})$ without privacy. These findings highlight how adaptive noise strategies can improve learning efficiency even under stringent privacy constraints.
@@
All R scripts used to generate the figures and reproduce the simulation results in this paper are openly available on our GitHub repository: https://github.com/YuxinLiu1997/Privacy-Preserving-Sequential-Learning.
@@ 6 Additional Related Work
$k=\lfloor\log_{\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}}\frac{1-p}{p}\rfloor+1.$ (14)
@@
$l_{3}=\log\frac{\mathbb{P}(x_{1}=+1|\theta=+1)}{\mathbb{P}(x_{1}=+1|\theta=-1)}\frac{\mathbb{P}(x_{2}=+1|\theta=+1,x_{1}=+1)}{\mathbb{P}(x_{2}=+1|\theta=-1,x_{1}=+1)}=\log\left[\frac{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}\right]^{2}.$
@@
$l_{3}=\log\frac{\mathbb{P}(x_{1}=-1|\theta=+1)}{\mathbb{P}(x_{1}=-1|\theta=-1)}\frac{\mathbb{P}(x_{2}=-1|\theta=+1,x_{1}=-1)}{\mathbb{P}(x_{2}=-1|\theta=+1,x_{1}=-1)}=\log\left[\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}\right]^{2}.$
@@
$\log\left[\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}\right]^{2}+\log\frac{p}{1-p}<0,\log\left[\frac{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}\right]^{2}+\log\frac{1-p}{p}>0.$
@@
In total, if $k$ is the threshold for the start of the information cascade, then $k=\lfloor\log_{\frac{(1-u(\varepsilon))(1-p)+u(\varepsilon)p}{u(\varepsilon)(1-p)+p(1-u(\varepsilon))}}\frac{1-p}{p}\rfloor+1$.
@@
When $\tilde{u}(\varepsilon)\neq\frac{1}{2}$, the probability of a correct cascade is given by:
$\mathbb{P}(\text{correct cascade})=\frac{\rho(\varepsilon)^{k}-1}{\rho(\varepsilon)^{2k}-1}.$
where $\rho(\varepsilon)=\frac{1-\tilde{u}(\varepsilon)}{\tilde{u}(\varepsilon)}$ and $\tilde{u}(\varepsilon)=u(\varepsilon)(1-p)+p(1-u(\varepsilon))$. When $\tilde{u}(\varepsilon)=\frac{1}{2}$, the information cascade does not occur.
@@
For the case $\tilde{u}(\varepsilon)=\frac{1}{2}$, it is trivial. However, for $\tilde{u}(\varepsilon)\neq\frac{1}{2}$, we can reframe the problem as a Markov chain featuring two absorbing states: $k$ and k-k. The transition probabilities are defined as follows: $\mathbb{P}(z_{n}=m+1|z_{n-1}=m)=\tilde{u}(\varepsilon)$ and $\mathbb{P}(z_{n}=m-1|z_{n-1}=m)=1-\tilde{u}(\varepsilon)$, where $z_{n-1}$ is the difference between the occurrences of action +1+1 and action 1-1 in the observation history up to agent $n$. This formulation aligns with the classical gambler’s ruin problem.
@@
$s_{2k}(k)=\tilde{u}(\varepsilon)s_{2k}(k+1)+(1-\tilde{u}(\varepsilon))s_{2k}(k-1).$
@@
$s_{2k}(k)=\frac{\rho(\varepsilon)^{k}-1}{\rho(\varepsilon)^{2k}-1}.$
where $\rho(\varepsilon)=\frac{1-\tilde{u}(\varepsilon)}{\tilde{u}(\varepsilon)}$ and $\tilde{u}(\varepsilon)=u(\varepsilon)(1-p)+p(1-u(\varepsilon))$. ∎
@@
where $\alpha=(1-p)/p$. Thus, $I_{k}=(\varepsilon_{k+1},\varepsilon_{k}]$, and $\varepsilon_{k}=\log\frac{1-v_{k}}{v_{k}}$. By lemma 8.2 take the first derivative of $s_{2k}(k)$ with respect to $\rho(\varepsilon)$ for $k\geq 2$ and $\rho(\varepsilon)\in(0,1)$:
$\frac{ds_{2k}(k)}{d\rho}=\frac{3k\rho(\varepsilon)^{3k-2}-k\rho(\varepsilon)^{k-1}-2k\rho(\varepsilon)^{2k-1}}{(\rho(\varepsilon)^{2k}-1)^{2}}<0.$
Next, we take the first derivative of $\rho(\varepsilon)$ and $u(\varepsilon)$ with respect to $u$ and $\varepsilon$, respectively. Then $\frac{d\rho(\varepsilon)}{du(\varepsilon)}>0$ and $\frac{du(\varepsilon)}{d\varepsilon}<0$. Consequently. $\frac{ds_{2k}(k)}{d\varepsilon}>0$, With an increase in the differential privacy budget, the probability of a correct cascade also increases.
9 Failure of Asymptotic Learning for $\varepsilon$-Differential Privacy
In this appendix we show that under an $\varepsilon$-differential privacy constraint, agents adopt a randomized response strategy with a fixed flip probability, and consequently asymptotic learning fails to occur in the sequential learning model with Gaussian signals.
Theorem 9.1 (Randomized Response Strategy for Continuous Model)
In the sequential learning model with Gaussian signals under an $\varepsilon$-differential privacy constraint, agents adopt a randomized response strategy with a fixed flip probability
$u_{n}=u(\varepsilon)=\frac{1}{1+e^{\varepsilon}},$
independent of the agent index $n$.
Proof.
Proof of Proposition 9.1 we can derive the decision for agent $n$ before adding any noise. A simple calculation shows that agent $n$ takes action +1+1 iff $l_{n}+L_{n}>0$ which implies
@@ 9 Failure of Asymptotic Learning for $\varepsilon$-Differential Privacy
$L_{n}=\log{\frac{\mathbb{P}(\theta=+1|s_{n})}{\mathbb{P}(\theta=-1|s_{n})}}=\log\frac{f(s_{n}|\theta=+1)}{f(s_{n}|\theta=-1)}=\log{\frac{e^{-(s_{n}-1)^{2}/2\sigma^{2}}}{e^{-(s_{n}+1)^{2}/2\sigma^{2}}}}=\frac{2s_{n}}{\sigma^{2}}.$
@@
where $\tilde{a_{n}}=-a_{n}$, then $u_{n}=\frac{1}{1+e^{\varepsilon}}$. ∎
Theorem 9.2 (Failure of Asymptotic Learning)
Proof.
Proof of Proposition 9.2 Without loss of generality, we assume $\theta=-1$, then the likelihood ratio of the public belief
@@
$=(1-u(\varepsilon))G_{-}(-l_{\infty})+u(\varepsilon)(1-G_{-}(-l_{\infty}))$
$=1-u(\varepsilon).$
In all, because agents always randomize actions for the purpose of protecting privacy with constant probability, asymptotic learning will not occur. ∎
@@
where $D_{+}(l_{n})=\log\frac{\mathbb{P}(x_{n}=+1|l_{n},\theta=+1)}{\mathbb{P}(x_{n}=+1|l_{n},\theta=-1)}$, $G_{-}(-l_{n})=a(e^{\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}}-e^{-\varepsilon+\frac{\varepsilon^{2}\sigma^{2}}{2}})e^{-\varepsilon\frac{\sigma^{2}l_{n}}{2}}$, and $a=\frac{1}{2}$.
@@
$\mathbb{P}_{\pi,-}(\sigma=+\infty)\geq\frac{4\alpha}{1+2\alpha}.$
@@
$\frac{\phi(\pi|x)}{1-\phi(\pi|x)}=\frac{\pi}{1-\pi}\cdot\frac{\mathbb{P}(x_{n}=-1|\pi,\theta=+1)}{\mathbb{P}(x_{n}=-1|\pi,\theta=-1)}.$
@@
$\mathbb{P}(x_{n}=-1|\pi,\theta=+1)<\mathbb{P}(x_{n}=-1|\pi,\theta=-1),$
@@
By Lemma 13.1, $\pi_{n+1}^{*}\geq\frac{1}{2}$, and by Lemma 13.3, we have
@@ 15 The Evolution of Log-Likelihood Ratio
Proof of Theorems 4.15 and 4.16: Learning Efficiency—Finite Expected Time to First Correct Action and Finite Expected Total Number of Incorrect Actions with Heterogeneous Privacy Budgets
Theorems 4.15 and 4.16 mirror the structure and argument of Theorems 4.10 and 4.11, which analyze learning efficiency in the homogeneous privacy setting. The key steps in the proof remain the same: by substituting the asymptotic learning rate result in Theorem 4.13, we obtain the corresponding expression of the expectations. Therefore, we omit the full proof and show only that the expectations in both theorems are finite.
@@ 19 Extension to Pufferfish Privacy
$u_{n}(s_{n})=\frac{e^{-k\varepsilon}}{1+e^{\varepsilon}},\quad\text{if }s_{n}\in(t(l_{n})+ka,\ t(l_{n})+(k+1)a)\cup(t(l_{n})-(k+1)a,\ t(l_{n})-ka),\ k\in\mathbb{N}.$
@@
When the parameter $a$ increases, if the ratio $\frac{a}{\varepsilon}$ is finite, then the asymptotic learning rate $f(n)=\frac{2a}{\varepsilon\sigma^{2}}\log(C_{\varepsilon,a}n)$ increases. This is because a larger $a$ induces a more asymmetric smooth randomized response, which improves the informativeness of the observed actions and thus accelerates the overall learning process. However, this improvement comes with a caveat: as $a\to\infty$, the term $C_{\varepsilon,a}$ tends to zero, which causes $f(n)$ to diverge. Therefore, we restrict our attention to the regime where $\frac{\varepsilon}{a}$ is constant and bounded away from zero to ensure that the learning rate remains well defined and informative.
@@
where $C_{1}$ is a positive constant that does not depend on $\varepsilon$ and $\zeta\left(x\right)=\sum_{n=1}^{\infty}n^{-x}$. If $\varepsilon<\frac{2a}{\sigma^{2}}$, the series converges and thus the expected time to the first correct action is finite:
@@
where $C_{2}$ is a positive constant that does not depend on $\varepsilon$. If and only if $\varepsilon<\frac{2a}{\sigma^{2}}$, the series converges and thus the expected total number of incorrect actions is finite:
@@
Although a higher value of $a$ makes it easier to ensure the finiteness of both $\mathbb{E}[\tau]$ and $\mathbb{E}[W]$, this benefit is not without potential drawbacks. In particular, as $a\to\infty$, the term $C_{\varepsilon,a}^{-\frac{2a}{\varepsilon\sigma^{2}}}$ diverges to infinity. Therefore, our results are valid only under the condition that $\frac{\varepsilon}{a}$ is constant and bounded away from zero. This ensures that the learning efficiency bounds remain meaningful and finite.